# Neural-Driven Image Editing

### v2

@@ A.4.2 Supplementary Case Studies
Qualitative comparisons are presented in Figures 10-13, corresponding to the four broad editing categories: Global, Background, Object, and Text Editing. For clarity and conciseness in the figures, the original lengthy instructions have been distilled into single-sentence descriptions without altering their intended meaning. The editing results of our neural-driven and neural-speech fusion methods consistently outperform text-prompt-based editing results, demonstrating superior alignment with human intent and greater editing precision. Notably, Text Editing presents a more complex challenge compared to other categories. Given the current limitations of backbone models (with the exception of commercial models like GPT-4o), text-based edits remain difficult. As evidenced by the examples, neural-driven approaches exhibit a stronger ability to align with human intent, making the editing process more intuitive and effective. It is foreseeable that in the near future, as reliable image-editing backbones become more accessible, neural-driven image editing will further stabilize and mature, evolving into an indispensable tool for everyday creative workflows.
@@
Figure 14 specifically analyzes three characteristic failure modes: (1) cases involving overly imaginative descriptions that deviate significantly from the training data distribution (e.g., "long-legged space creature"), (2) ambiguous instructions with insufficient semantic details (particularly evident in case (b) where background retention specifications were omitted), and (3) challenges posed by non-standard input image dimensions (such as panoramic aspect ratios). These failure cases provide valuable insights into the current limitations of neural-based editing systems.
@@
Figure 14: Qualitative analysis on three failure cases: (a) Overly exaggerated descriptions, e.g., "long-legged space creature"; (b) Vague instructions lacking detail, such as omitting whether to retain the background; (c) Uncommon image dimensions, e.g., panoramic input images.
A.5 Limitations Discussion
A.6 Broader Societal Impacts
A.7 Safety and Ethics

### v3

@@ E.3 More Qualitative Results
More qualitative comparisons are presented in Figures 10-13, corresponding to the four broad editing categories: Global, Background, Object, and Text Editing. For clarity and conciseness in the figures, the original lengthy instructions have been distilled into single-sentence descriptions without altering their intended meaning. The editing results of our neural-driven and neural-speech fusion methods consistently outperform text-prompt-based editing results, demonstrating superior alignment with human intent and greater editing precision. Notably, Text Editing presents a more complex challenge compared to other categories. Given the current limitations of backbone models (with the exception of commercial models like GPT-4o), text-based edits remain difficult. As evidenced by the examples, neural-driven approaches exhibit a stronger ability to align with human intent, making the editing process more intuitive and effective. It is foreseeable that in the near future, as reliable image-editing backbones become more accessible, neural-driven image editing will further stabilize and mature, evolving into an indispensable tool for everyday creative workflows.
@@
Fig. 14 specifically analyzes three characteristic failure modes: (1) cases involving overly imaginative descriptions that deviate significantly from the training data distribution (e.g., "long-legged space creature"), (2) ambiguous instructions with insufficient semantic details (particularly evident in case (b) where background retention specifications were omitted), and (3) challenges posed by non-standard input image dimensions (such as panoramic aspect ratios). These failure cases provide valuable insights into the current limitations of neural-based editing systems.
E.4 More Failure Cases
Figure. 14 specifically illustrates three failure cases, where overly exaggerated imagination, vague instructions, or uncommon input image sizes may contribute to failed results. It demonstrates three characteristic failure modes: (1) cases involving overly imaginative descriptions that deviate significantly from the training data distribution (e.g., "long-legged space creature"), (2) ambiguous instructions with insufficient semantic details (particularly evident in case (b) where background retention specifications were omitted), and (3) challenges posed by non-standard input image dimensions (such as panoramic aspect ratios). These failure cases provide valuable insights into the current limitations of neural-based editing systems.
@@ E.4 More Failure Cases
Figure 14: Qualitative analysiscomparison onf our neural-driven and speech-neural fusion methods and text-prompt baseline for three failure cases: (a) Overly exaggerated descriptions, e.g., "long-legged space creature"; (b) Vague instructions lacking detail, such as omitting whether to retain the background; (c) Uncommon image dimensions, e.g., panoramic input images.